# Conditional Entropies of

### v1

@@ I Introduction
In the last decade, channels that introduce insertion and deletion errors attracted significant attention due to their relevance to DNA storage systems [4, 33, 50, 51, 72, 76], where deletions and insertions are among the most dominant errors [38, 59]. The study of communication channels with insertion and deletion errors is also relevant to many other applications such as the synchronization of files and symbols of data streams [24] and for cases of over-sampling and under-sampling at the receiver side [60]. VT codes, designed by Varshamov and Tenengolts [73], are the first family of codes that can correct a single deletion or a single insertion [44]. Later, several works extended the scheme to correct multiple deletion errors, as well as substitutions and edit errors; see e.g., [14, 15, 28, 30, 34, 63, 18, 64, 65, 62, 70, 71]. Under some applications, such as DNA storage systems, the problem of list decoding was studied as well, for example in [3, 35, 37, 40, 47, 74]. Under this setup, the decoder receives a channel output and returns a (short) list of possible codewords that includes the transmitted codeword. Even though many works considered channels with deletion errors and designed codes that correct such errors, finding optimal codes for these channels is yet far from being solved.
@@
The rest of the paper is organized as follows. In Section II, we motivate the task of studying the input and output entropies of the $k$-deletion and $k$-insertion channels by describing its relation to the problem of analysing the capacity of deletion channels. In Section III, we give the basic notations and definitions in the paper. In this section, we also give a formal definition of the problems studied in this paper. In Section IV, we characterize the input entropies of the $k$-deletion and $k$-insertion channels and particularly for $k=1$ $2$, we give a more elaborate characterization. Section V finds the extremum values and the sequences that attain these extremum values for the input entropy of both $1$-deletion and $1$-insertion channels over arbitrary alphabet size. Section VI studies the average input entropies for $k=1$. In Section VII and Section VIII, we find the minimum values and the sequences that attain these minimum values for the input entropy of $2$-deletion channel and $k$-insertion channel, respectively, over binary alphabet.
@@ II Motivation
For a given input distribution $\mathsf{Pr}({\boldsymbol{x}})$ over $X_{1}^{n}$, the mutual information $I(X_{1}^{n};Y)$, where $Y=k\textrm{-}\mathsf{Del}(X_{1}^{n})$, is calculated by
@@
$=H(Y)-\sum_{{\boldsymbol{x}}\in\{0,1\}^{n}}\mathsf{Pr}({\boldsymbol{x}})H(Y|X_{1}^{n}={\boldsymbol{x}}).$
Or, alternatively,
$I(X^{n}_{1};Y)$ $=H(X_{1}^{n})-H(X_{1}^{n}~{}|~{}Y)$
$=H(X_{1}^{n})-\sum_{{\boldsymbol{y}}\in\{0,1\}^{n-k}}\mathsf{Pr}({\boldsymbol{y}})H(X^{n}_{1}|Y={\boldsymbol{y}}).$
Hence, one of the important tasks in studying the capacity of synchronization channels is to determine for every ${\boldsymbol{x}}\in{\cal X}^{n}$ the conditional entropy $H(Y~{}|~{}X^{n}_{1}={\boldsymbol{x}})$ and similarly for every ${\boldsymbol{y}}\in{\cal Y}^{*}$ the conditional entropy $H(X^{n}_{1}~{}|~{}Y={\boldsymbol{y}})$. Formally, for an input ${\boldsymbol{x}}\in{\cal X}^{n}$, we refer to $H(Y~{}|~{}X^{n}_{1}={\boldsymbol{x}})$ as the output entropy of the channel and for an output ${\boldsymbol{y}}\in{\cal Y}^{*}$, $H(X^{n}_{1}~{}|~{}Y={\boldsymbol{y}})$ is the input entropy of the channel.
@@
Among other applications, the output entropy can be used in list-decoders to rank the different codewords from the decoder’s list by their transmission probability [35, 74]. The output entropy can also be used to assist with code design by prioritizing sequences with lower output entropy since they are likely to have higher successful decoding probability. Alternatively, when the goal is to encrypt the information, sequences with higher output entropies are preferred [7]. In several of DNA reconstruction algorithms such as [10, 57, 69], there is a limitation on the number of noisy copies that can considered by the algorithm’s decoder, due to design restrictions and run-time considerations. Therefore, in case the number of received noisy copies is greater than this limitation, a subset of these copies should be considered. Hence, to improve the accuracy of such algorithms, the input entropy of the channel outputs can be used to wisely select this subset of copies. Studying the input entropy was also investigated in [6, 7] for analyzing the information leakage of a key by revealing any of its subsequences.
@@ III Preliminaries and Problem Statement
Let $\mathbb{I}$ denote the indicator function. For a positive integer $n$, let $[n]\triangleq\{1,\ldots,n\}$ and let $\Sigma_{q}\triangleq\{0,1,\ldots,q-1\}$ be an alphabet of size $q$. For an integer $n\geq 0$, let $\Sigma_{q}^{n}$ be the set of all words (sequences) of length $n$ over the alphabet $\Sigma_{q}$. For an integer $k$, $0\leq k\leq n$, a sequence ${\boldsymbol{y}}\in\Sigma_{q}^{n-k}$ is a $k$-subsequence of ${\boldsymbol{x}}\in\Sigma_{q}^{n}$ if ${\boldsymbol{y}}$ can be obtained by deleting $k$ symbols from ${\boldsymbol{x}}$. Similarly, a sequence ${\boldsymbol{y}}\in\Sigma_{q}^{n+k}$ is a $k$-supersequence of ${\boldsymbol{x}}\in\Sigma_{q}^{n}$ if ${\boldsymbol{x}}$ is a $k$-subsequence of ${\boldsymbol{y}}$. Let ${\boldsymbol{x}}$ and ${\boldsymbol{y}}$ be two sequences of length $n$ and $m$ respectively such that $m<n$. The embedding number of ${\boldsymbol{y}}$ in ${\boldsymbol{x}}$, denoted by $\omega_{{\boldsymbol{y}}}({\boldsymbol{x}})$, is defined as the number of distinct occurrences of ${\boldsymbol{y}}$ as a subsequence of ${\boldsymbol{x}}$. More formally, the embedding number is the number of distinct index sets, $(i_{1},i_{2},\ldots,i_{m})$ , such that $1\leq i_{1}<i_{2}<\cdots<i_{m}\leq n$ and $x_{i_{1}}=y_{1},x_{i_{2}}=y_{2},\dotsc,x_{i_{m}}=y_{m}$. For example, for ${\boldsymbol{x}}=\texttt{11220}$ and ${\boldsymbol{y}}=\texttt{120}$, it holds $\omega_{{\boldsymbol{y}}}({\boldsymbol{x}})=4$. The $k$-insertion ball centred at ${{\boldsymbol{x}}\in\Sigma_{q}^{n}}$, denoted by $I_{k}({\boldsymbol{x}})\subseteq\Sigma_{q}^{n+k}$, is the set of all $k$-supersequences of ${\boldsymbol{x}}$. Similarly, the $k$-deletion ball centred at ${{\boldsymbol{x}}\in\Sigma_{q}^{n}}$, denoted by $D_{k}({\boldsymbol{x}})\subseteq\Sigma_{q}^{n-k}$, is the set of all $k$-subsequences of ${\boldsymbol{x}}$.
@@
$\mathsf{Pr}_{k\textrm{-}\mathsf{Del}}^{\mathsf{Out}}\{{\boldsymbol{y}}~{}|~{}{\boldsymbol{x}}\}=\frac{\omega_{{\boldsymbol{y}}}({\boldsymbol{x}})}{\binom{n}{k}},$
@@
$\mathsf{Pr}_{k\textrm{-}\mathsf{Ins}}^{\mathsf{Out}}\{{\boldsymbol{y}}~{}|~{}{\boldsymbol{x}}\}=\frac{\omega_{{\boldsymbol{x}}}({\boldsymbol{y}})}{\binom{n+k}{k}q^{k}},$
@@
    For all ${\boldsymbol{x}}\in\Sigma_{q}^{n}$, find its output entropy over the channel $k\textrm{-}\mathsf{Del}$,
    $\mathsf{H}^{\mathsf{Out}}_{k\textrm{-}\mathsf{Del}}({\boldsymbol{x}})\triangleq H(Y~{}|~{}X^{n}_{1}={\boldsymbol{x}}).$
@@
    Find the minimum, maximum, and average output entropy of the channel $k\textrm{-}\mathsf{Del}$,
@@
    For all ${\boldsymbol{y}}\in\Sigma_{q}^{n-k}$, find its input entropy over the channel $k\textrm{-}\mathsf{Del}$ with input distribution $\mathsf{Pr}(X^{n}_{1})$,
    $\mathsf{H}^{\mathsf{In}}_{k\textrm{-}\mathsf{Del},\mathsf{Pr}(X^{n}_{1})}({\boldsymbol{y}})\triangleq H(X^{n}_{1}~{}|~{}Y={\boldsymbol{y}}).$
@@
    Find the minimum, maximum, and average input entropy of the channel $k\textrm{-}\mathsf{Del}$ with input distribution $\mathsf{Pr}(X^{n}_{1})$,
@@
Atashpendar et al. [6] studied the minimum input entropy of the $k\textrm{-}\mathsf{Del}$ channel for the case where $k\in\{1,2\}$ and $q=2$. In their work, they presented an efficient algorithm to count the number of occurrences of a sequence ${\boldsymbol{y}}$ in any given supersequence ${\boldsymbol{x}}$. They also provided an algorithm that receives a sequence ${\boldsymbol{y}}$, computes the set of all of its supersequences of specific length, characterizes the distribution of their embedding numbers and clusters them by their Hamming weight. Lastly, they proved that the all-zero and the all-one sequences minimize the input entropy of the $1\textrm{-}\mathsf{Del}$ and the $2\textrm{-}\mathsf{Del}$ channels, under uniform transmission.
@@
Based on Lemma 1, the rest of the paper studies the input entropy of the $k$-insertion and the $k$-deletion channels. Unless otherwise stated, we shall assume uniform transmission and thus, we drop the $\mathsf{Pr}(X^{n}_{1})$ term from the notation of the input entropy.
@@ IV Characterization of the Input Entropy
Proof.
$\mathsf{H}^{\mathsf{In}}_{k\textrm{-}\mathsf{Del}}({\boldsymbol{y}})$ $=-\sum_{{\boldsymbol{x}}\in I_{k}({\boldsymbol{y}})}\left(\mathsf{Pr}_{k\textrm{-}\mathsf{Del}}^{\mathsf{In}}\{{\boldsymbol{x}}~{}|~{}{\boldsymbol{y}}\}\cdot\log\left(\mathsf{Pr}_{k\textrm{-}\mathsf{Del}}^{\mathsf{In}}\{{\boldsymbol{x}}~{}|~{}{\boldsymbol{y}}\}\right)\right)$
@@
Proof.
$\mathsf{H}^{\mathsf{In}}_{k\textrm{-}\mathsf{Ins}}({\boldsymbol{y}})$ $=-\sum_{{\boldsymbol{x}}\in D_{k}({\boldsymbol{y}})}\left(\mathsf{Pr}_{k\textrm{-}\mathsf{Ins}}^{\mathsf{In}}\{{\boldsymbol{x}}~{}|~{}{\boldsymbol{y}}\}\cdot\log\left(\mathsf{Pr}_{k\textrm{-}\mathsf{Ins}}^{\mathsf{In}}\{{\boldsymbol{x}}~{}|~{}{\boldsymbol{y}}\}\right)\right)$
@@
For a sequence ${\boldsymbol{x}}$, a run of ${\boldsymbol{x}}$ is a maximal subsequence of identical consecutive symbols within ${\boldsymbol{x}}$. The number of runs in ${\boldsymbol{x}}$ is denoted by $\rho({\boldsymbol{x}})$. We denote by $\Sigma_{q,R}^{n}$, the set of sequences ${\boldsymbol{x}}\in\Sigma_{q}^{n}$, such that $\rho({\boldsymbol{x}})=R$. It is well known that the number of such sequences is ${}|~{}\Sigma_{q,R}^{n}~{}|~{}={\binom{n-1}{R-1}}q(q-1)^{R-1}$. For a sequence ${\boldsymbol{x}}\in\Sigma_{q,R}^{n}$, its run length profile, denoted by ${\cal R}{\cal L}({\boldsymbol{x}})$, is the vector of the lengths of the runs in ${\boldsymbol{x}}$. That is, ${\cal R}{\cal L}({\boldsymbol{x}})\triangleq(r_{1},r_{2},\ldots,r_{R}),$ where $r_{i}$, for $i\in[R]$ denotes the length of the $i$-th run of ${\boldsymbol{x}}$. For example, for $q=4$ and ${\boldsymbol{x}}=\texttt{311221110}$, we have that ${\cal R}{\cal L}({\boldsymbol{x}})=(1,2,2,3,1)$.
For the special case of $k=1$, we have the next two theorems for the $1\textrm{-}\mathsf{Del}$ and $1\textrm{-}\mathsf{Ins}$ channels.
@@
Proof.
From [11, 56], $I_{1}({\boldsymbol{y}})$ has $\rho({\boldsymbol{y}})$ sequences that can be obtained by prolonging an existing run, each with embedding number $r_{i}+1$ for $1\leq i\leq\rho({\boldsymbol{y}})$. The embedding number of the remaining ${}|~{}I_{1}({\boldsymbol{y}})~{}|~{}-\rho({\boldsymbol{y}})=q+~{}|~{}{\boldsymbol{y}}~{}|~{}(q-1)-\rho({\boldsymbol{y}})$ sequences is one. Hence, upon substituting these embedding numbers in the expression from Lemma 2, we get
@@
Proof.
From [11, 56], $D_{1}({\boldsymbol{y}})$ has $\rho({\boldsymbol{y}})$ sequences with embedding numbers, $r_{1},r_{2},\ldots,r_{\rho({\boldsymbol{y}})}$, obtained by shortening a run. Hence, upon substituting these embedding numbers in the expression from Lemma 3, we get
@@
For the special case of $k=2$, we have the next two theorems for the $2\textrm{-}\mathsf{Del}$ and $2\textrm{-}\mathsf{Ins}$ channels.
@@
V Extremum Values for $k=1$

### v2

@@ I Introduction
In the last decade, channels that introduce insertion and deletion errors attracted significant attention due to their relevance to DNA storage systems [3, 35, 52, 53, 75, 79], where deletions and insertions are among the most dominant errors [40, 61]. The study of communication channels with insertion and deletion errors is also relevant to many other applications such as the synchronization of files and symbols of data streams [25] and for cases of over-sampling and under-sampling at the receiver side [62]. VT codes, designed by Varshamov and Tenengolts [76], are the first family of codes that can correct a single deletion or a single insertion [46]. Later, several works extended the scheme to correct multiple deletion errors, as well as substitutions and edit errors; see e.g., [14, 15, 30, 32, 36, 65, 18, 66, 67, 64, 73, 74, 22, 71]. Under some applications, such as DNA storage systems, the problem of list decoding was studied as well, for example in [2, 37, 39, 42, 49, 77]. Under this setup, the decoder receives a channel output and returns a (short) list of possible codewords that includes the transmitted codeword. Even though many works considered channels with deletion errors and designed codes that correct such errors, finding optimal codes for these channels is yet far from being solved.
@@
The rest of the paper is organized as follows. In Section II, we motivate the task of studying the input and output entropies of the $k$-deletion and $k$-insertion channels by describing its relation to the problem of analysing the capacity of deletion channels. In Section III, we give the basic notations and definitions in the paper. In this section, we also give a formal definition of the problems studied in this paper. In Section IV, we characterize the input entropies of the $k$-deletion and $k$-insertion channels and particularly for $k=1$ $2$, we give a more elaborate characterization. Section V finds the extremum values and the sequences that attain these extremum values for the input entropy of both $1$-deletion and $1$-insertion channels over arbitrary alphabet size. Section V also studies the average input entropies for $k=1$. In Section VI, we find the minimum values and the sequences that attain these minimum values for the input entropy of $2$-deletion channel and $k$-insertion channel over binary alphabet. Section VII then concludes the paper with a brief summary of our contribution and discussion on future work and open problems.
@@ II Motivation
For a given input distribution $\mathsf{Pr}({\boldsymbol{x}})$ over $X_{1}^{n}$, the mutual information $I(X_{1}^{n};Y)$, where $Y=k\textrm{-}\mathsf{Del}(X_{1}^{n})$, is given by
@@
$=H(Y)-\sum_{{\boldsymbol{x}}\in\{0,1\}^{n}}\mathsf{Pr}({\boldsymbol{x}})\,H(Y|X_{1}^{n}={\boldsymbol{x}}),$
or, equivalently,
$I(X_{1}^{n};Y)$ $=H(X_{1}^{n})-H(X_{1}^{n}|Y)$
$=H(X_{1}^{n})-\sum_{{\boldsymbol{y}}\in\{0,1\}^{n-k}}\mathsf{Pr}({\boldsymbol{y}})\,H(X_{1}^{n}|Y={\boldsymbol{y}}).$
Hence, one of the important tasks in studying the capacity of synchronization channels is to determine for every ${\boldsymbol{x}}\in{\cal X}^{n}$ the conditional entropy $H(Y|X^{n}_{1}={\boldsymbol{x}})$ and similarly for every ${\boldsymbol{y}}\in{\cal Y}^{*}$ the conditional entropy $H(X^{n}_{1}|Y={\boldsymbol{y}})$. Formally, for an input ${\boldsymbol{x}}\in{\cal X}^{n}$, we refer to $H(Y|X^{n}_{1}={\boldsymbol{x}})$ as the output entropy of the channel and for an output ${\boldsymbol{y}}\in{\cal Y}^{*}$, $H(X^{n}_{1}|Y={\boldsymbol{y}})$ is the input entropy of the channel.
@@
Remark 1.
In this paper, we study the conditional entropies under uniform transmission. Nevertheless, a complete characterization of the channel capacity necessitates a deeper understanding of the optimal input distribution, which remains an important open problem and is left for future investigation.
Among other applications, the output entropy can be used in list-decoders to rank the different codewords from the decoder’s list by their transmission probability [37, 77]. The output entropy can also be used to assist with code design by prioritizing sequences with lower output entropy since they are likely to have higher successful decoding probability. Alternatively, when the goal is to encrypt the information, sequences with higher output entropies are preferred [5]. In several of DNA reconstruction algorithms such as [10, 59, 72], there is a limitation on the number of noisy copies that can be considered by the algorithm’s decoder, due to design restrictions and run-time considerations. Therefore, in case the number of received noisy copies is greater than this limitation, a subset of these copies should be considered. Hence, to improve the accuracy of such algorithms, the input entropy of the channel outputs can be used to wisely select this subset of copies. Studying the input entropy was also investigated in [6, 5] for analyzing the information leakage of a key by revealing any of its subsequences.
@@ III Preliminaries and Problem Statement
Let $\mathbb{I}$ denote the indicator function. For a positive integer $n$, let $[n]\triangleq\{1,\ldots,n\}$ and let $\Sigma_{q}\triangleq\{0,1,\ldots,q-1\}$ be an alphabet of size $q$. For an integer $n\geq 0$, let $\Sigma_{q}^{n}$ be the set of all words (sequences) of length $n$ over the alphabet $\Sigma_{q}$. For an integer $k$, $0\leq k\leq n$, a sequence ${\boldsymbol{y}}\in\Sigma_{q}^{n-k}$ is a $k$-subsequence of ${\boldsymbol{x}}\in\Sigma_{q}^{n}$ if ${\boldsymbol{y}}$ can be obtained by deleting $k$ symbols from ${\boldsymbol{x}}$. Similarly, a sequence ${\boldsymbol{y}}\in\Sigma_{q}^{n+k}$ is a $k$-supersequence of ${\boldsymbol{x}}\in\Sigma_{q}^{n}$ if ${\boldsymbol{x}}$ is a $k$-subsequence of ${\boldsymbol{y}}$. Let ${\boldsymbol{x}}$ and ${\boldsymbol{y}}$ be two sequences of length $n$ and $m$ respectively such that $m<n$. The embedding number of ${\boldsymbol{y}}$ in ${\boldsymbol{x}}$, denoted by $\omega_{{\boldsymbol{y}}}({\boldsymbol{x}})$, is defined as the number of distinct occurrences of ${\boldsymbol{y}}$ as a subsequence of ${\boldsymbol{x}}$. More formally, the embedding number is the number of distinct index sets, $(i_{1},i_{2},\ldots,i_{m})$, such that $1\leq i_{1}<i_{2}<\cdots<i_{m}\leq n$ and $x_{i_{1}}=y_{1},x_{i_{2}}=y_{2},\dotsc,x_{i_{m}}=y_{m}$. For example, let ${\boldsymbol{x}}=\texttt{11220}$ and ${\boldsymbol{y}}=\texttt{120}$, then $\omega_{{\boldsymbol{y}}}({\boldsymbol{x}})=4$. The $k$-insertion ball centred at ${{\boldsymbol{x}}\in\Sigma_{q}^{n}}$, denoted by $I_{k}({\boldsymbol{x}})\subseteq\Sigma_{q}^{n+k}$, is the set of all $k$-supersequences of ${\boldsymbol{x}}$. Similarly, the $k$-deletion ball centred at ${{\boldsymbol{x}}\in\Sigma_{q}^{n}}$, denoted by $D_{k}({\boldsymbol{x}})\subseteq\Sigma_{q}^{n-k}$, is the set of all $k$-subsequences of ${\boldsymbol{x}}$.
@@
$\mathsf{Pr}_{k\textrm{-}\mathsf{Del}}^{\mathsf{Out}}\{{\boldsymbol{y}}|{\boldsymbol{x}}\}=\frac{\omega_{{\boldsymbol{y}}}({\boldsymbol{x}})}{\binom{n}{k}},$
@@
$\mathsf{Pr}_{k\textrm{-}\mathsf{Ins}}^{\mathsf{Out}}\{{\boldsymbol{y}}|{\boldsymbol{x}}\}=\frac{\omega_{{\boldsymbol{x}}}({\boldsymbol{y}})}{\binom{n+k}{k}q^{k}},$
@@
    For all ${\boldsymbol{x}}\in\Sigma_{q}^{n}$, find its output entropy,
    $\mathsf{H}^{\mathsf{Out}}_{k\textrm{-}\mathsf{Del}}({\boldsymbol{x}})\triangleq H(Y|X^{n}_{1}={\boldsymbol{x}}).$
@@
    Find the minimum, maximum, and average output entropy,
@@
    For all ${\boldsymbol{y}}\in\Sigma_{q}^{n-k}$, find its input entropy with input distribution $\mathsf{Pr}(X^{n}_{1})$,
    $\mathsf{H}^{\mathsf{In}}_{k\textrm{-}\mathsf{Del},\mathsf{Pr}(X^{n}_{1})}({\boldsymbol{y}})\triangleq H(X^{n}_{1}|Y={\boldsymbol{y}}).$
@@
    Find the minimum, maximum, and average input entropy with input distribution $\mathsf{Pr}(X^{n}_{1})$,
@@
Remark 2.
Building on Lemma 1, the remainder of the paper analyzes the input entropy of the $k$-insertion and $k$-deletion channels under the uniform input distribution. Accordingly, we omit the $\mathsf{Pr}(X_{1}^{n})$ term from the notation for input entropy throughout the rest of the paper.
III-A Related Work
Atashpendar et al. [6] studied the minimum input entropy of the $k\textrm{-}\mathsf{Del}$ channel under uniform transmission for the case where $k\in\{1,2\}$ and $q=2$. In their work, they presented an efficient algorithm to count the number of occurrences of a sequence ${\boldsymbol{y}}$ in any given supersequence ${\boldsymbol{x}}$. They also provided an algorithm that receives a sequence ${\boldsymbol{y}}$, computes the set of all of its supersequences of specific length, characterizes the distribution of their embedding numbers and clusters them by their Hamming weight. Lastly, they proved that the all-zero and the all-one sequences minimize the input entropy of the $1\textrm{-}\mathsf{Del}$ and the $2\textrm{-}\mathsf{Del}$ channels.
In another related line of work, Atashpendar et al. [7] studied the minimum input entropy of the $k\textrm{-}\mathsf{Del}$ channel under uniform transmission within an asymptotic regime where the output length is fixed while the input length and the parameter $k$ tend to infinity. Their approach employs analytic-combinatorial techniques based on the hidden word statistics framework developed by Flajolet et al.[29]. In particular, they analyze the moments of the posterior distribution (i.e., the conditional probability of an input sequence given a fixed output sequence) and demonstrate that these are governed by an input-dependent autocorrelation measure. They show that the all-zero and all-one sequences maximize this autocorrelation measure, which in turn leads to the minimization of the input entropy.
III-B Our Contributions
This paper makes the following key contributions:
  1. 1.
    General entropy characterizations. We derive explicit expressions for the input and output conditional entropies of the $k$-deletion and $k$-insertion channels under the assumption of uniform input distribution, expressing them in terms of embedding numbers of sequences.
  2. 2.
    Closed-form formulas for $k=1,2$.
    • For the 1-deletion and 1-insertion channels, we obtain entropy formulas invariant under permutations of the run-length profile, reducing the analysis to combinatorial properties of runs.
    • For the 2-deletion channel, we extend prior work of Atashpendar et al. [6] by providing new closed-form entropy expressions that incorporate the structure of alternating segments.
  3. 3.
    Extremal entropy values and extremizing sequences.
    • We fully characterize the sequences that attain the minimum and maximum input entropy in both $1$-deletion and $1$-insertion channels: single-run sequences minimize input entropy, while alternating sequences maximize it. In particular, we resolve the conjecture of Atashpendar et al. that alternating sequences maximize the input entropy for $1$-deletion channel.
    • We prove that single-run sequences uniquely minimize the input entropy for $2$-deletion channel and determine the exact minimum value. In doing so, we improve upon the previous characterization in the literature, employ a refined alternating-segment analysis, and rigorously establish both the minimizing sequences and the exact minimum input entropy value for the $2$-deletion channel.
  4. 4.
    Average entropy We derive exact formulas for the average input entropy of $1$-insertion and $1$-deletion channels by enumerating the contribution of each possible run length across all sequences, and we establish closed-form lower bounds.
  5. 5.
    Results for higher-order channels. For arbitrary $k\geq 2$, we prove that constant sequences uniquely minimize the input entropy for the $k$-insertion channel.
@@ IV Characterization of the Input Entropy
Proof:
$\mathsf{H}^{\mathsf{In}}_{k\textrm{-}\mathsf{Del}}({\boldsymbol{y}})$ $=-\sum_{{\boldsymbol{x}}\in I_{k}({\boldsymbol{y}})}\left(\mathsf{Pr}_{k\textrm{-}\mathsf{Del}}^{\mathsf{In}}\{{\boldsymbol{x}}|{\boldsymbol{y}}\}\cdot\log\left(\mathsf{Pr}_{k\textrm{-}\mathsf{Del}}^{\mathsf{In}}\{{\boldsymbol{x}}|{\boldsymbol{y}}\}\right)\right)$
@@
Proof:
$\mathsf{H}^{\mathsf{In}}_{k\textrm{-}\mathsf{Ins}}({\boldsymbol{y}})$ $=-\sum_{{\boldsymbol{x}}\in D_{k}({\boldsymbol{y}})}\left(\mathsf{Pr}_{k\textrm{-}\mathsf{Ins}}^{\mathsf{In}}\{{\boldsymbol{x}}|{\boldsymbol{y}}\}\cdot\log\left(\mathsf{Pr}_{k\textrm{-}\mathsf{Ins}}^{\mathsf{In}}\{{\boldsymbol{x}}|{\boldsymbol{y}}\}\right)\right)$
@@
For a sequence ${\boldsymbol{x}}$, a run of ${\boldsymbol{x}}$ is a maximal subsequence of identical consecutive symbols within ${\boldsymbol{x}}$. The number of runs in ${\boldsymbol{x}}$ is denoted by $\rho({\boldsymbol{x}})$. We denote by $\Sigma_{q,R}^{n}$, the set of sequences ${\boldsymbol{x}}\in\Sigma_{q}^{n}$, such that $\rho({\boldsymbol{x}})=R$. It is well known that the number of such sequences is $|\Sigma_{q,R}^{n}|={\binom{n-1}{R-1}}q(q-1)^{R-1}$. For a sequence ${\boldsymbol{x}}\in\Sigma_{q,R}^{n}$, its run length profile, denoted by ${\cal R}{\cal L}({\boldsymbol{x}})$, is the vector of the lengths of the runs in ${\boldsymbol{x}}$. That is, ${\cal R}{\cal L}({\boldsymbol{x}})\triangleq(r_{1},r_{2},\ldots,r_{R}),$ where $r_{i}$, for $i\in[R]$, denotes the length of the $i$-th run of ${\boldsymbol{x}}$. For example, let $q=4$ and ${\boldsymbol{x}}=\texttt{311221110}$, then ${\cal R}{\cal L}({\boldsymbol{x}})=(1,2,2,3,1)$.
For $k=1$, we have the next two theorems for the $1\textrm{-}\mathsf{Del}$ and $1\textrm{-}\mathsf{Ins}$ channels.
@@
Proof:
From [11, 58], it is known that $I_{1}({\boldsymbol{y}})$ contains $\rho({\boldsymbol{y}})$ sequences that can be obtained by prolonging an existing run, each with embedding number $r_{i}+1$ for $1\leq i\leq\rho({\boldsymbol{y}})$. The embedding number of the remaining $|I_{1}({\boldsymbol{y}})|-\rho({\boldsymbol{y}})=q+|{\boldsymbol{y}}|(q-1)-\rho({\boldsymbol{y}})$ sequences is one. Hence, upon substituting these embedding numbers in the expression from Lemma 2, we get
@@
Proof:
From [11, 58], it is known that $D_{1}({\boldsymbol{y}})$ contains $\rho({\boldsymbol{y}})$ sequences with embedding numbers, $r_{1},r_{2},\ldots,r_{\rho({\boldsymbol{y}})}$, obtained by shortening a run. Hence, upon substituting these embedding numbers in the expression from Lemma 3, we get
@@
For $k=2$, we have the next two theorems for the $2\textrm{-}\mathsf{Del}$ and $2\textrm{-}\mathsf{Ins}$ channels.
@@
The result then follows by substituting these embedding numbers and their respective counts into the input entropy expression derived in Lemma 2. ∎
@@
The result then follows by substituting these embedding numbers and their respective counts into the input entropy expression derived in Lemma 3. ∎
V Extremum and Average Values for $k=1$